OpenStudy (anonymous):

plz help me: Q1) In slope-intercept form, y=mx+b where y=?m=?x=?b=?

OpenStudy (hba):

Slope Intercept Form = y = mx + b where b = y-intercept of the line, m = Slope of the line

OpenStudy (anonymous):

@Hero @him1618 @ash2326 help me

OpenStudy (hba):

@Hares333 Got it ?

OpenStudy (anonymous):

what do you mean? @hba

OpenStudy (anonymous):

b = y-intercept of the line,?????

OpenStudy (hba):

Yes

OpenStudy (anonymous):

b=y how?

OpenStudy (hba):

y-intercept is actually the point on the graph where x is zero

OpenStudy (anonymous):

ok!! so what's x

OpenStudy (anonymous):

@jiteshmeghwal9

OpenStudy (hba):

x is x

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

can you explain me with the help of diagram @hba

OpenStudy (ash2326):

@Hares333 y=mx+b This equation represents a line, with slope m and y-intercept b x and y are variables, There are infinite points on a line, these represents the points' coordinates. For each unique x there is a unique y.  The point where the line intersects the y-axis is the y intercept and obviously x being 0, if you put x=0. y=0x+b=b

OpenStudy (anonymous):

thnx... @ash2326

OpenStudy (ash2326):

Welcome, did you understand? @Hares333

OpenStudy (anonymous):

yaa @ash2326